a)Price of Pepsi is Rs.120 and its quantity demanded is 300 bottles. Suppose due to Raman Pepsi company reduces price of it to Rs. 108 ,due to it its demand increases to 360 bottles. Calculate elasticity and also mention the type of elasticity.
1
Expert's answer
2020-04-29T09:21:45-0400

Price elasticity of demand is equal to change in quantity demand of a good as result of change in price of the good.

Ped= "\frac{change in quanity demanded}{change in price}"

Change in quantity demanded = "\frac{Q_i-Q_o}{(Q_o+Q_i)/2}\times100"

="\frac{360-300}{(360+300)/2}\times100"

=18.18

Change in price = "\frac{P_i-P_o}{(P_i+P_o)/2}\times100"

="\frac{108-120}{(108+120)/2}\times100"

=-8.33

Ped= "\frac{18.18}{-8.33}"

Ped= -2.18

Since the price elasticity of demand is greater than 1 in absolute value, the demand for Pepsi is elastic